solve by substitution method
0.1x+0.3y=-0.7
0.9x-0.2y=8.2

Answer by stanbon(75887)

solve by substitution method
0.1x+0.3y=-0.7
0.9x-0.2y=8.2
----
1st: Get rid of all those decimals
x + 3y = -7
9x - 2y = 82
-----
Multiply the 1st equation by "9" to get:
9x + 27y = -63
9x - 2y = 82
----
Subtract and solvd for "y":
29y = 145
y = 5
----
Solve for "x":
x + 3y = -7
x + 15 = -7
x = -22/3
================
cheers,
Stan H.
=============